OpenStudy (zubhanwc3):

if dy/dx = cos^2 (piy/4) and if y = 1 when x = 0, then when y = 3, x is equal to

OpenStudy (zubhanwc3):

if \[\frac{ dy }{ dx } = \cos ^{2}(\frac{ \pi y }{ 4 })\]and if y = 1 when x = 0, then when y = 3, x is equal to 1/8 -pi/8 1 8/-pi 8/pi

ganeshie8 (ganeshie8):

use double angle identity : \(\cos (2 x) = 2\cos^2x - 1 \)

ganeshie8 (ganeshie8):

and solve the diff eq'n first

ganeshie8 (ganeshie8):

\[\frac{ dy }{ dx } = \cos ^{2}(\frac{ \pi y }{ 4 })\] \[\frac{ dy }{ dx } = \dfrac{1 + \cos( 2 \frac{\pi y}{4})}{2}\]

ganeshie8 (ganeshie8):

integrate